OpenStudy (anonymous):

how to find factor of 8400

OpenStudy (aum):

prime factorization?

OpenStudy (aum):

8400 = 2 * 4200 = 2 * 2 * 2100 = 2 * 2 * 2 * 1050 = 2 * 2 * 2 * 2 * 525 = 2 * 2 * 2 * 2 * 5 * 105 = .....?....

OpenStudy (wolf1728):

You could use a prime factor calculator http://1728.org/prime.htm and your answer will be 2*2*2*2*3*5*5*7

OpenStudy (wolf1728):

and here's what Wikipedia says about finding prime factors: http://en.wikipedia.org/wiki/Prime_factor